OpenStudy (neo92):

solve \[\int\limits_{}^{} \sqrt{a^{2}-x^{2}} \] using integration by parts

OpenStudy (neo92):

\[\int\limits_{}^{} \sqrt{a^{2}-x^{2}} dx\]

OpenStudy (lgbasallote):

it's specific that you use integration by parts?

OpenStudy (lgbasallote):

or is that just your preference?

OpenStudy (neo92):

yep.. i knw it can b done easily using a substitution. but gotta do it by parts.. :(

OpenStudy (lgbasallote):

well you have no choice...you use \[ u = \sqrt{a^2 - x^2}\] \[dv = dx\]

OpenStudy (lgbasallote):

\[du = -\frac{x}{\sqrt{a^2 - x^2}}\] \[v = x\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (lgbasallote):

\[uv - \int vdu\] \[\implies x\sqrt{a^2 - x^2} + \int \frac{x^2}{\sqrt{a^2-x^2}}dx\] now that is integrable easily right?

OpenStudy (lgbasallote):

do i still need to proceed? or do you know what to do now?

OpenStudy (neo92):

wait i'l try n tel ya

OpenStudy (lgbasallote):

sure. go for it

OpenStudy (lgbasallote):

i stand corrected by my earier words...\(\int \frac{x^2}{\sqrt{a^2 - x^2}}dx\) is NOT an easy integral =))

OpenStudy (lgbasallote):

you're gonna have to use trig sub one way or another...

OpenStudy (anonymous):

This is a fairly evil integral (look up the answer in a table), and as you say, integration by parts is not going to get you anywhere. I'm not sure I can work it out myself without reference to Strang or some other source. Do you know Strang's book? It's online.

OpenStudy (anonymous):

Or Paul's notes must have it too.

OpenStudy (lgbasallote):

actuallly that last integral i presented \[\int \frac{x^2}{\sqrt{a^2-x^2}}dx\] is solveable...but it involves integration by parts + trig sub

OpenStudy (lgbasallote):

so it's still possible to use IPB

OpenStudy (anonymous):

Maybe so, looking at the answer, but it'll be several steps.

OpenStudy (neo92):

@lgbasallote yep.. i too agree with u.

OpenStudy (lgbasallote):

but anyway...yes it is evil

OpenStudy (mimi_x3):

not that evil :) try \(x=asin\theta\)

OpenStudy (mimi_x3):

then simplify; not that hard i suppose

OpenStudy (neo92):

@Mimi_x3 using a substitution i okay.. but my bro said his teacher asked him 2 do it usin IBP.. thats da prob

OpenStudy (mimi_x3):

didnt igba use IBP then reduced it to that integral? that is what i can see above; from that trig sub and you're done

OpenStudy (mimi_x3):

im talking about this integral \[\int\limits\frac{x^{2}}{\sqrt{a^{2}-x^{2}}} dx\]

OpenStudy (neo92):

yep.. I think theres noway to solve \int \frac{x^2}{\sqrt{a^2 - x^2}}dx without using trig sub? isnt it?

OpenStudy (neo92):

\[\int\limits \frac{x^2}{\sqrt{a^2 - x^2}}dx\]

OpenStudy (mimi_x3):

well if you want to do integration by parts twice; i think it is possible i havent tried tho. but why bother? since you already did integration by parts once? just trig sub for the last step

OpenStudy (neo92):

itz nt me whois bothering.. itz ma brother.. :D thnk all of ya for da help.